# Distill Any Depth: Distillation Creates a Stronger Monocular Depth Estimator

### v1

@@ Abstract
Monocular depth estimation (MDE) aims to predict scene depth from a single RGB image and plays a crucial role in 3D scene understanding. Recent advances in zero-shot MDE leverage normalized depth representations and distillation-based learning to improve generalization across diverse scenes. However, current depth normalization methods for distillation, relying on global normalization, can amplify noisy pseudo-labels, reducing distillation effectiveness. In this paper, we systematically analyze the impact of different depth normalization strategies on pseudo-label distillation. Based on our findings, we propose Cross-Context Distillation, which integrates global and local depth cues to enhance pseudo-label quality. Additionally, we introduce a multi-teacher distillation framework that leverages complementary strengths of different depth estimation models, leading to more robust and accurate depth predictions. Extensive experiments on benchmark datasets demonstrate that our approach significantly outperforms state-of-the-art methods, both quantitatively and qualitatively.
@@ 1 Introduction
Monocular depth estimation (MDE) predicts scene depth from a single RGB image, offering flexibility compared to stereo or multi-view methods. This makes MDE ideal for applications like autonomous driving and robotic navigation [10, 12, 16, 48, 28]. Recent research on zero-shot MDE models [34, 51, 43, 22] aims to handle diverse scenarios, but training such models requires large-scale, diverse depth data, which is often limited by the need for specialized equipment [29, 50]. A promising solution is using large-scale unlabeled data, which has shown success in tasks like classification and segmentation [25, 58, 44]. Studies like DepthAnything [46] highlight the effectiveness of using pseudo labels from teacher models for training student models.
@@
Building on this insight, in this paper, we first investigate the issue of depth normalization in pseudo-label distillation. We begin by analyzing various depth normalization strategies, including global normalization, local normalization, hybrid global-local approaches, and the absence of normalization. Through empirical experiments, we explore how each technique affects the performance of various distillation designs, especially when using pseudo-labels for training. Our analysis provides valuable insights into how different normalization methods influence the MDE loss function and distillation outcomes, offering a set of best practices for optimizing performance in diverse scenarios.
Building on this empirical foundation, we introduce a Cross-Context Distillation method, designed to distill knowledge from the teacher model more effectively. We are motivated by our finding that local regions, when used for distillation, produce pseudo-labels that capture higher-quality depth details, improving the student model’s depth estimation accuracy. However, focusing solely on local regions might overlook the broader contextual relationships in the image. To address the issue , we combine local and global inputs within a unified distillation framework. By combining the context-specific advantages of local distillation with the broader understanding provided by global methods, our method achieves more detailed and reliable depth predictions.
Furthermore, we propose a multi-teacher distillation framework that leverages the complementary strengths of multiple depth estimation models. Our design is motivated by the observation of recent advancements that diffusion-based models, benefiting from large-scale image priors, excel at capturing fine-grained details but are computationally expensive, while encoder-decoder models provide higher accuracy and efficiency but relatively lack fine-detail reconstruction. To harness these strengths, we randomly select different models to generate pseudo-labels, and then supervise the student model based on these labels. This operation enables the student model to learn from the detailed depth information of diffusion-based models while benefiting from the precision of encoder-decoder models.
To validate the effectiveness of our design, we conduct extensive experiments on various benchmark datasets. The empirical results show that our method significantly outperforms existing baselines qualitatively and quantitatively. The contributions can be summarized below:
  • We systematically analyze the role of different depth normalization strategies in pseudo-label distillation, providing insights into their effects on MDE performance.
  • We propose Cross-Context Distillation, a hybrid local-global distillation framework that enhances distillation by leveraging both fine-grained details and global depth relationships.
  • We develop a multi-teacher distillation framework that integrates pseudo labels from multiple depth estimation models, combining the strengths of various depth models.
  • We conduct extensive experiments on benchmark datasets, demonstrating that our method outperforms state-of-the-art approaches both quantitatively and qualitatively. Code and models are made publicly available.
@@ 2.1 Monocular Depth Estimation
Monocular depth estimation (MDE) has evolved from hand-crafted methods to deep learning, significantly improving accuracy [10, 27, 11, 15, 57, 35]. Architectural refinements, such as multi-scale designs and attention mechanisms, have further enhanced feature extraction [19, 5, 56]. However, most models remain reliant on labeled data and struggle to generalize across diverse environments. Zero-shot MDE improves generalization by leveraging large-scale datasets, geometric constraints, and multi-task learning [34, 51, 53, 55]. Metric depth estimation incorporates intrinsic data for absolute depth learning [2, 52, 20, 42], while generative models such as Marigold refine depth details using diffusion priors [22, 45]. Despite these advances, effectively utilizing unlabeled data remains a challenge due to pseudo-label noise and inconsistencies across different contexts. DepthAnything [47] explores large-scale unlabeled data but struggles with pseudo-label reliability. PatchFusion [8, 30] improves depth estimation by refining high-resolution image representations but lacks adaptability in generative settings. To address these issues, we propose Cross-Context and Multi-Teacher Distillation, which enhances pseudo-label supervision by leveraging diverse contextual information and multiple expert models, improving both accuracy and generalization ability.
@@ 2.2 Semi-supervised Monocular Depth Estimation
Semi-supervised depth estimation has gained attention by utilizing temporal consistency to better use unlabeled data [26, 17]. Some methods [1, 40, 6, 49, 14] apply stereo geometric constraints, enforcing left-right consistency to enhance depth accuracy, while others use additional supervision like semantic priors [33, 18] or GANs, such as DepthGAN [21]. However, these approaches are limited by their reliance on temporal cues or stereo constraints, restricting their applicability. Recent work [32] explored pseudo-labeling for semi-supervised MDE but lacks generative modeling capabilities. DepthAnything [46] demonstrated the potential of large-scale unlabeled data, though pseudo-label reliability remains challenging. In contrast, our approach improves pseudo-label reliability and enhances MDE accuracy, relying solely on unlabeled data without additional constraints.
@@ 3 Method
In this section, we introduce a novel distillation framework designed to leverage unlabeled images for training zero-shot Monocular Depth Estimation (MDE) models. We begin by exploring various depth normalization techniques in Section 3.1, followed by detailing our proposed distillation method in Section 3.2, which combines predictions across multiple contexts. The overall framework is illustrated in Fig. 3. Finally, we describe a multi-teacher distillation mechanism in Section 3.2 that integrates diverse depth estimators as teacher models to train the student model.
@@ 3.2 Distillation Pipeline
In this section, we introduce an enhanced distillation pipeline that integrates two complementary strategies: Cross-Context Distillation and Multi-Teacher Distillation. Both strategies aim to improve the quality of pseudo-label distillation, enhance the model’s fine-grained perception, and boost generalization across diverse scenarios.
@@
Cross-context Distillation.
A key challenge in monocular depth distillation is the trade-off between local detail preservation and global depth consistency. As shown in Fig. 5, providing a local crop of an image as input to the teacher model enhances fine-grained details in the pseudo-depth labels, but it may fail to capture the overall scene structure. Conversely, using the entire image as input preserves the global depth structure but often lacks fine details. To address this limitation, we propose Cross-Context Distillation, a method that enables the student model to learn both local details and global structures simultaneously. Cross-context distillation consists of two key strategies:
@@
Figure 6: Multi-teacher Mechanism. We introduce a multi-teacher distillation approach, where pseudo-labels are generated from multiple teacher models. At each training iteration, one teacher is randomly selected to produce pseudo-labels for unlabeled images.
Multi-teacher Distillation.
In addition to cross-context distillation, we adopt a multi-teacher distillation strategy, illustrated in Fig. 6, to further enhance the quality and robustness of the distilled depth knowledge. This approach leverages multiple teacher models, each trained with distinct architectures, optimization strategies, or data distributions, to generate diverse pseudo-labels. By aggregating knowledge from multiple sources, the student model benefits from a richer and more generalized depth representation. Formally, given a set of pre-trained teacher models $\mathcal{M}_{1},\mathcal{M}_{2},\dots,\mathcal{M}_{N}$, we employ a probabilistic teacher selection mechanism, where one teacher model is randomly selected at each training iteration to generate pseudo-labels for the input image. The inclusion of multiple teacher models allows the student to learn from a diverse set of predictions, effectively mitigating biases and limitations inherent to any single model.
@@ 4.1 Experimental Settings
Datasets. To evaluate the effectiveness of our proposed distillation framework, we follow the methodology outlined in DepthAnythingv2 [47]. Specifically, we conduct our distillation experiments using a subset of 200,000 samples from the SA-1B dataset [24].
For evaluation, we assess the performance of the distilled student model on five widely used depth estimation benchmarks, ensuring that these datasets remain unseen during training to enable a robust zero-shot evaluation. The chosen benchmarks include: NYUv2 [39], KITTI [13], ETH3D [38], ScanNet [7], and DIODE [41]. Additional dataset details are provided in the Appendix.
@@
Implementation. Our experiments use state-of-the-art monocular depth estimation models as teachers to generate pseudo-labels, supervising various student models in a distillation framework with only RGB images as input. In shared-context distillation, both teacher and student receive the same global region, extracted via random cropping from the original image. The crop maintains a 1:1 aspect ratio and is sampled within a range from 644 pixels to the shortest side of the image, then resized to $560\times 560$ for local predictions. In global-local distillation, the global region is cropped into overlapping local patches, each sized $560\times 560$, for the teacher model to predict pseudo-labels. We use GenPercept [45]) and DepthAnythingv2 (DAv2) [47] as teacher models for the multi-teacher mechanism. The learning rate is in tune with that of the corresponding student model. For DAv2 [47], the decoder learning rate is set to $5\times 10^{-5}$. For the total loss function, we set the parameters as follows: $\lambda_{1}=0.5$, $\lambda_{2}=1.0$ and $\lambda_{3}=2.0$.
@@ 4.2 Analysis
For the ablation study and analysis, we sample a subset of 50K images from SA-1B [24] as our training data, with an input image size of 560 × 560 for the network. We conduct experiments on two of the most challenging benchmarks, DIODE [41] and ETH3D [38], which include both indoor and outdoor scenes. The model was trained with a batch size of 4 and converged after approximately 20,000 iterations on a single NVIDIA V100 GPU.
Impact of Normalization across Cross-Context Distillation. We evaluate the effect of different normalization strategies on Cross-Context Distillation, as shown in Table 1. The results indicate that the optimal normalization method varies across different distillation strategies. For shared-context distillation, no normalization achieves the best performance, assuming that pseudo-depth labels naturally reside in the same domain as predictions. For Local-Global distillation, Hybrid Normalization proves most effective, maintaining consistent depth predictions across regions through hierarchical normalization within specific depth ranges.
@@
Ablation Study of Cross-Context Distillation. To further validate the effectiveness of our distillation framework, we conduct ablation studies by removing Shared-Context Distillation and Local-Global Distillation in Table 2. The results show that both components contribute significantly to
@@
Shared-Context Distillation Global Norm. $0.067$ $0.243$
No Norm. 0.058 0.236
Local Norm. $0.060$ $0.238$
Hybrid Norm. $0.059$ $0.237$
Local-Global Distillation Global Norm. $0.065$ $0.253$
No Norm. $0.060$ $0.235$
Local Norm. $0.059$ $0.235$
Hybrid Norm. 0.056 0.232
@@
improving the student model’s ability to utilize pseudo-labels, demonstrating the robustness of our approach.
@@
Cross-Architecture Distillation. To evaluate our normalization strategy, we conducted experiments using MiDaS [34] and DepthAnything [47], testing four configurations (DA-L, MiDaS-L, DA-S, MiDaS-S) as shown in Table 3. Our method consistently outperforms previous distillation approaches that use global normalization on the DIODE [41] and ETH3D [38] datasets, demonstrating superior performance both within and across architectures, and highlighting the limitations of global normalization in pseudo-label distillation.
Multi-teacher Mechanism. We evaluate the effectiveness of our multi-teacher distillation strategy across five benchmarks in Table 4. To handle the diverse output depth distributions of different teacher models, we use Hybrid Normalization for Shared-Context Distillation in this experiment. Using diffusion-based Genpercept [45] and Dinov2-based DepthAnythingv2 [47] as teacher models, we train a lightweight DPT-based depth estimation model. Our approach outperforms both teacher models overall, with only a minor gap on KITTI [13], demonstrating the effectiveness of multi-teacher distillation.
Table 4: Quantitative comparison of our multi-teacher distillation model on zero-shot benchmarks. The bold values indicate the best performance. Our model, which integrates diverse depth estimation models, achieves higher accuracy than any individual teacher model.
Method NYUv2 KITTI DIODE ScanNet ETH3D Avg. Rank
DepthAnything v2 NeurIPS’24 0.045 0.979 0.074 0.946 0.262 0.754 0.042 0.978 0.131 0.865 1.9
Genpercept(Disparity)ICLR’25 0.058 0.969 0.080 0.934 0.226 0.741 0.063 0.960 0.096 0.959 2.6
Ours(Multi-teacher) 0.043 0.981 0.077 0.945 0.298 0.756 0.042 0.979 0.065 0.983 1.4
@@
Ours 0.043 0.981 0.070 0.949 0.233 0.753 0.042 0.980 0.054 0.981
  • Refers to our method applied on the MiDaS v3.1. Refers to our method applied on the DepthAnythingv2-Large.

### v2

@@ Abstract
Recent advances in zero-shot monocular depth estimation(MDE) have significantly improved generalization by unifying depth distributions through normalized depth representations and by leveraging large-scale unlabeled data via pseudo-label distillation. However, existing methods that rely on global depth normalization treat all depth values equally, which can amplify noise in pseudo-labels and reduce distillation effectiveness. In this paper, we present a systematic analysis of depth normalization strategies in the context of pseudo-label distillation. Our study shows that, under recent distillation paradigms (e.g., shared-context distillation), normalization is not always necessary—omitting it can help mitigate the impact of noisy supervision. Furthermore, rather than focusing solely on how depth information is represented, we propose Cross-Context Distillation, which integrates both global and local depth cues to enhance pseudo-label quality. We also introduce an assistant-guided distillation strategy that incorporates complementary depth priors from a diffusion-based teacher model, enhancing supervision diversity and robustness. Extensive experiments on benchmark datasets demonstrate that our approach significantly outperforms state-of-the-art methods, both quantitatively and qualitatively.
@@ 1 Introduction
Monocular depth estimation (MDE) predicts scene depth from a single RGB image, offering greater flexibility compared to stereo or multi-view methods, and benefiting a wide range of applications, such as autonomous driving and robotic perception [11, 13, 18, 52, 30]. Recent research on zero-shot MDE models [37, 55, 47, 24] aims to handle diverse scenarios, but training such models requires large-scale, diverse depth data, which is often limited by the need for specialized equipment [31, 54]. A promising solution is using large-scale unlabeled data, which has shown success in tasks like classification and segmentation [27, 63, 48]. Studies like DepthAnything [50] highlight the effectiveness of using pseudo labels from teacher models for training student models.
@@
Building on this insight, in this paper, we investigate the issue of depth normalization in pseudo-label distillation. We analyze various depth normalization strategies, including global normalization, local normalization, hybrid global-local approaches, and the absence of normalization. Through empirical experiments, we explore how each depth representation impacts the performance of different distillation designs, especially when using pseudo-labels for training.
Rather than focusing solely on pseudo-label representation, we introduce Cross-Context Distillation, a method that integrates both global and local depth cues to enhance pseudo-label quality. Our findings reveal that local regions, when used for distillation, produce pseudo-labels that capture higher-quality depth details, improving the student model’s depth estimation accuracy. However, relying solely on local regions may overlook broader contextual relationships in the image. To address this, we combine both local and global inputs within a unified distillation framework. By leveraging the context-specific advantages of local distillation alongside the broader understanding provided by global methods, our approach yields more detailed and reliable depth predictions.
To harness the strengths of both, we propose using a diffusion-based model as the teacher assistant to generate pseudo-labels, which are then used to supervise the student model. This strategy enables the student model to learn from the detailed depth information provided by diffusion-based models, while also benefiting from the precision and efficiency of encoder-decoder models.
To validate the effectiveness of our design, we conduct extensive experiments on various benchmark datasets. The empirical results show that our method significantly outperforms existing baselines qualitatively and quantitatively. The contributions can be summarized below: 1) We systematically analyze the role of different depth normalization strategies in pseudo-label distillation, providing insights into their effects on MDE performance. 2) To enhance the quality of pseudo-labels, we propose Cross-Context Distillation, a hybrid local-global framework that leverages fine-grained details and global depth relationships; a teacher assistant that harnesses the complementary strengths of diverse depth estimation models to further improve robustness and accuracy. 3) We conduct extensive experiments on benchmark datasets, demonstrating that our method outperforms state-of-the-art approaches both quantitatively and qualitatively.
@@ 2.1 Monocular Depth Estimation
Monocular depth estimation (MDE) has evolved from hand-crafted methods to deep learning, significantly improving accuracy [11, 29, 12, 16, 62, 38]. Architectural refinements, such as multi-scale designs and attention mechanisms, have further enhanced feature extraction [21, 6, 61]. However, most models remain reliant on labeled data and struggle to generalize across diverse environments. Zero-shot MDE improves generalization by leveraging large-scale datasets, geometric constraints, and multi-task learning [37, 55, 57, 60, 58]. Metric depth estimation incorporates intrinsic data for absolute depth learning [2, 56, 22, 35, 46, 4], while generative models such as Marigold refine depth details using diffusion priors [24, 49, 17, he2024lotus]. Despite these advances, effectively utilizing unlabeled data remains a challenge due to pseudo-label noise and inconsistencies across different contexts. DepthAnything [51] explores large-scale unlabeled data but struggles with pseudo-label reliability. PatchFusion [9, 32] improves depth estimation by refining high-resolution image representations but lacks adaptability in generative settings. To address these issues, we propose Cross-Context and Multi-Teacher Distillation, which enhances pseudo-label supervision by leveraging diverse contextual information and multiple expert models, improving both accuracy and generalization ability.
@@ 2.2 Semi-supervised Monocular Depth Estimation
Semi-supervised depth estimation has garnered increasing attention, primarily leveraging temporal consistency to utilize unlabeled data more effectively [28, 19]. Some approaches [1, 44, 7, 53, 15] integrate stereo geometric constraints, enforcing left-right consistency in stereo video to enhance depth accuracy. Others incorporate additional supervision, such as semantic priors [36, 20]or generative adversarial networks (GANs). For instance, DepthGAN [23] refines depth predictions through adversarial learning. However, these methods often rely on temporal cues, stereo constraints, or other auxiliary information, limiting their applicability to broader and more general scenarios. Recent work [34] has explored pseudo-labeling for semi-supervised monocular depth estimation (MDE), but it lacks generative modeling capabilities, restricting its generalization across diverse environments. DepthAnything  [50] demonstrates the effectiveness of large-scale unlabeled data in improving generalization; however, pseudo-label reliability remains a challenge. In contrast, our approach focuses on single-image depth estimation, improving pseudo-label reliability and maximizing its effectiveness. By relying solely on unlabeled data without additional constraints, our method achieves a more accurate and generalizable MDE model.
@@ 3 Method
In this section, we introduce a novel distillation framework designed to leverage unlabeled images for training zero-shot Monocular Depth Estimation (MDE) models. We begin by exploring various depth normalization techniques in Section 3.1, followed by detailing our proposed distillation method in Section 3.2, which combines predictions across multiple contexts. The overall framework is illustrated in Fig. 3. Finally, we introduce an assistant-guided distillation scheme, in which a diffusion-based model acts as an auxiliary teacher to provide additional supervision for student training.
@@ 3.2 Distillation Pipeline
In this section, we introduce an enhanced distillation pipeline that integrates two complementary strategies: Cross-Context Distillation andassistant-guided distillation. Both strategies aim to improve the quality of pseudo-label distillation, enhance the model’s fine-grained perception.
Cross-context Distillation. A key challenge in monocular depth distillation is the trade-off between local detail preservation and global depth consistency. As shown in Fig. 5, providing a local crop of an image as input to the teacher model enhances fine-grained details in the pseudo-depth labels, but it may fail to capture the overall scene structure. Conversely, using the entire image as input preserves the global depth structure but often lacks fine details. To address this limitation, we propose Cross-Context Distillation, a method that enables the student model to learn both local details and global structures simultaneously. Cross-context distillation consists of two key strategies:
@@
Assistant-Guided Distillation. In addition to cross-context distillation, we propose an assistant-guided distillation strategy to further enhance the quality and robustness of the distilled depth knowledge. This approach pairs a primary teacher [51] with a single auxiliary assistant, selected as a diffusion-based depth estimator [49], which leverages generative priors to complement the primary teacher’s predictions. This design leverages their complementary strengths: the primary teacher excels in providing efficient and globally consistent supervision, while the assistant offers fine-grained depth cues derived from its generative modeling capabilities. By drawing supervision from two distinct architectures trained with different paradigms(e.g., optimization strategies or data distributions), the student benefits from diverse knowledge sources, effectively mitigating biases and limitations inherent to a single teacher model. Formally, let $\mathcal{M}$ and $\mathcal{M}_{a}$ denote the primary and assistant models, respectively. During training, a probabilistic sampling mechanism determines whether supervision for each iteration is drawn from $\mathcal{M}$ or $\mathcal{M}_{a}$. This stochastic guidance encourages the student to adapt to multiple supervision styles, fostering richer and more generalizable depth representations. Overall, this assistant-guided scheme introduces complementary and diversified pseudo-labels, reducing over-reliance on any single teacher and improving both generalization and depth estimation performance.
@@ 4.1 Experimental Settings
Datasets. We train our proposed distillation framework on a subset of 200,000 unlabeled images from the SA-1B dataset [26], following the training protocol of DepthAnythingv2 [51]. For evaluation, we assess the distilled student model on five widely used depth estimation benchmarks. All test datasets are kept unseen during training, enabling a zero-shot evaluation of generalization performance. The chosen benchmarks include: NYUv2 [43], KITTI [14], ETH3D [42], ScanNet [8], and DIODE [45]. Additional dataset details are provided in the Appendix.
@@
Implementation. Our experiments use state-of-the-art monocular depth estimation models as teachers to generate pseudo-labels, supervising various student models in a distillation framework with only RGB images as input. In shared-context distillation, both teacher and student receive the same global region, extracted via random cropping from the original image. The crop maintains a 1:1 aspect ratio and is sampled within a range from 644 pixels to the shortest side of the image, then resized to $560\times 560$ for prediction. In global-local distillation, the global region is cropped into overlapping local patches, each sized $560\times 560$, for the teacher model to predict pseudo-labels. For assistant-guided distillation, we adopt a probabilistic sampling strategy where the primary teacher and the assistant model are selected with a ratio of 7:3, respectively. We train our best student model using the distillation pipeline for 20,000 iterations with a batch size of 8 on a single NVIDIA V100 GPU, initialized with pre-trained DAv2-Large weights. The learning rate is in tune with that of the corresponding student model. For DAv2 [51], the decoder learning rate is set to $5\times 10^{-5}$. For the total loss function, we set the parameters as follows: $\lambda_{1}=0.5$, $\lambda_{2}=1.0$ and $\lambda_{3}=2.0$.
@@ 4.2 Analysis
For the ablation study and analysis, we sample a subset of 50K images from SA-1B [26] as our training data, with an input image size of 560 × 560 for the network. We conduct experiments on two of the most challenging benchmarks, DIODE [45] and ETH3D [42], which include both indoor and outdoor scenes.
Analysis of Normalization across Cross-Context Distillation. We evaluate the impact of different depth normalization strategies on Cross-Context Distillation, as shown in Table 1. The results reveal that the optimal normalization method depends on the specific distillation design. In shared-context distillation, where all pseudo-labels are generated by a single teacher model, hybrid normalization achieves the best performance, closely followed by no normalization. The consistent domain across supervision signals reduces the need for normalization, enabling the model to better preserve local depth relationships. Unlike ground-truth-based training—where normalization is essential to align depth distributions across datasets captured by heterogeneous sensors or represented in varying formats (e.g., sparse vs. dense, relative vs. absolute)—pseudo-labels from a single model are inherently more uniform. Therefore, direct L1 loss without normalization can more faithfully supervise pixel-level depth without distortion from global rescaling. In contrast, global normalization introduces undesirable inter-pixel dependencies, while local normalization discards global structural coherence. In local-global distillation, hybrid normalization again proves most effective, likely due to its hierarchical design that enforces consistency across both local and global depth predictions. The relatively small gap between hybrid and global normalization suggests that our framework, which uses local cues to refine global predictions, effectively mitigates the limitations of global normalization. However, no normalization leads to a notable performance drop compared to the shared-context setting, indicating that localized regions in this case come from distinct depth domains, making direct L1 supervision less reliable. Local normalization, as before, sacrifices global consistency and thus underperforms.
@@
Shared-Context Distillation Global Norm. $0.064$ $0.259$
No Norm. 0.057 $0.239$
Local Norm. $0.070$ $0.245$
Hybrid Norm. 0.057 0.238
Local-Global Distillation Global Norm. $0.065$ $0.239$
No Norm. $0.273$ $0.300$
Local Norm. $0.076$ $0.244$
Hybrid Norm. 0.064 0.238
@@
Ablation Study of Cross-Context Distillation. To further validate the effectiveness of our distillation framework, we conduct ablation studies by removing Shared-Context Distillation and Local-Global Distillation in Table 2. Without both components, the model degrades to a conventional distillation setup, resulting in significantly lower performance. Introducing Shared-Context Distillation with Hybrid Normalization notably improves accuracy, highlighting the benefits of a better normalization strategy with consistent context supervision. When using only Local-Global Distillation, the model still performs well, showing the effectiveness of region-wise depth refinement even without global context information. Combining both strategies yields the best results, confirming that both components contribute significantly to improving the student model’s ability to utilize pseudo-labels, demonstrating the robustness of our approach.
Cross-Architecture Distillation. To highlight the limitations of previous state-of-the-art distillation approaches employing global normalization, we compare their performance against the Hybrid Normalization strategy, which we utilize in our distillation framework, across diverse model architectures. To demonstrate the generalizability of our approach, we conduct cross-architecture distillation experiments on both the state-of-the-art DepthAnything [51] and the classic MiDaS [37] architecture. Experiments are conducted using MiDaS [37] and DepthAnything [51] in four configurations (DA-L, MiDaS-L, DA-S, MiDaS-S), as shown in Table 3. Our method consistently outperforms previous global normalization-based distillation on both the DIODE [45] and ETH3D [42] datasets. These results demonstrate superior performance both within and across architectures, underscoring the limitations of global normalization in pseudo-label distillation.
Table 4: Effect of Assistant-Guided Distillation. Bold values indicate the best performance. Our method integrates a primary teacher, DepthAnything v2 (denoted as ‘D’), with a diffusion-based assistant, GenPercept (denoted as ‘G’), leveraging their complementary strengths to produce higher-quality pseudo-labels. The student model, trained under this assistant-guided distillation framework, consistently achieves better accuracy than when distilled from the DAv2-Large teacher alone.
Method Assistant-Guided Strategy ETH3D DIODE
AbsRel↓ AbsRel↓
DepthAnything v2 w/o 0.131 0.262
Genpercept(Disparity) w/o 0.096 0.226
D + G Avg. 0.228 0.371
D + G Select. 0.054 0.258
Effect of Assistant-Guided Distillation. To validate the effectiveness of our proposedassistant-guided distillation strategy, we design a comparative experiment that introduces an additional assistant model to the conventional teacher-student distillation framework. Specifically, we adopt DepthAnything v2 as the primary teacher and GenPercept—a diffusion-based model—as the assistant. The student model shares the same architecture as DAv2-Large and is initialized with its pre-trained weights. This setup allows us to investigate whether supervision from two diverse architectures—trained under different paradigms—can offer complementary guidance that enhances both generalization and depth estimation performance. To explore the most effective way to combine pseudo-labels from the primary teacher and the assistant, we compare two assistant-guided strategies: (1) a weighted averaging approach (Avg.), which assigns greater weight to pixels where the two teachers exhibit high agreement, and (2) a selection-based strategy (Select.), which probabilistically samples the supervision signal from either teacher. While the averaging strategy attempts to leverage consistency between teachers, it often performs poorly due to conflicting pseudo-labels, where averaging can amplify errors. In contrast, the selection-based strategy allows the student to selectively absorb the strengths of each teacher, avoiding error reinforcement. As shown in Table 4, the Select. strategy significantly outperforms both individual teachers and the averaging method on the ETH3D benchmark, demonstrating the effectiveness ofassistant-guided distillation in delivering robust and diverse supervision.
@@
Ours 0.043 0.981 0.070 0.949 0.233 0.753 0.043 0.980 0.054 0.981
  • Cross-Context distillation applied to MiDaS v3.1, using a pre-trained MiDaS v3.1 model as the teacher.
  • Cross-Context distillation applied to DepthAnythingv2-Large, using a pre-trained DAv2 model as the teacher.